use hardy wineberg equation to find equilbrium.

AA 150. Aa 120. aa 90
Answer #1

The steps involved here

Step 1: Calculation number of alleles

Genotype

Freequency

Allele A

Allele a

Total

AA

150

300

0

300

Aa

120

120

120

240

aa

90

0

180

180

Total

360

420

300

720

Step 2: Calculation of allele frequencies

Frequency = No of a alleles/Total no of alleles

Allele A

= 420/720

= 0.58

Allele a

= 300/720

= 0.42

Step 3: Expected genotype frequencies estimation

AA

= 0.58*0.58= 0.340

*360= 123

Aa

= 2*0.58* 0.42= 0.486

*360= 175

aa

0.42*0.42= 0.174

*360= 63

Step 4: Chi-square test

Null hypothesis: The observed values are not deviating from the expected values.

Category

AA

Aa

aa

Observed values

150

120

90

Exprected Values

123

175

63

Deviation

28

-55

28

D^2

756.25

3025

756.25

D^2/E

6.17

17.29

12.10

35.56

X^2

35.56

Degrees of freedom

2

Inference: The calculated chisquar value i.e. 35.56 is greater than the table value i.e.5.99 at 2 DF and 0.05 probability, hence the null hypothesis is rejected. Which means the population in not in HW equilibrium.